# Multi: Multimodal Understanding Leaderboard with Text and Images

### v2

@@ Abstract
Rapid progress in multimodal large language models (MLLMs) highlights the need to introduce challenging yet realistic benchmarks to the academic community, while existing benchmarks primarily focus on understanding simple natural images and short context. In this paper, we present Multi as a cutting-edge benchmark for evaluating MLLMs on understanding complex tables and images and reasoning with long context. Multi provides multimodal inputs and requires responses that are either precise or open-ended, reflecting real-life examination styles. Multi includes over 18,000 questions, and challenges MLLMs with a variety of tasks, ranging from formula derivation to image detail analysis and cross-modality reasoning. We also introduce Multi-Elite, a 500-question selected hard subset, and Multi-Extend with more than 4,500 external knowledge context pieces. Our evaluation indicates significant potential for MLLM advancement, with GPT-4V achieving a 63.7% accuracy rate on Multi, in contrast to other MLLMs scoring between 28.5% and 55.3%. Multi serves not only as a robust evaluation platform but also paves the way for the development of expert-level AI. Details and access are available at: https://OpenDFM.github.io/MULTI-Benchmark/.
@@
Zichen Zhu and Yang Xu and Lu Chen22footnotemark: 2and Jingkai Yang and Yichuan Ma Yiming Sun and Hailin Wen and Jiaqi Liu and Jinyu Cai Yingzi Ma and Situo Zhang and Zihan Zhao and Liangtai Sun and Kai Yu22footnotemark: 2 X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China {JamesZhutheThird, xuyang0112, chenlusz, kai.yu}@sjtu.edu.cn
@@
Figure 2: The overview of Multi.
1 Introduction
The rapid advancement in large-scale language models (LLMs) has led to significant achievements in natural language processing and related disciplines. Yet, human communication and understanding extend beyond language, encompassing images, tables, mathematical and chemical formulas, graphs, diagrams, cartoons, posters, and other visual mediums. They play a crucial role in conveying information, particularly in scientific areas. Therefore, there’s a growing interest in developing Multimodal LLMs (MLLMs) capable of processing and generating across various modalities, including visual ones, and performing tasks that require cross-modal reasoning.
Evaluating MLLMs presents unique challenges. Current benchmarks Lu et al. (2022); Li et al. (2023b); Yue et al. (2023) either focus narrowly on natural scene images or are simplistic, failing to thoroughly assess the models’ abilities. Many scientific benchmarks Sun et al. (2023a); Huang et al. (2023) rely on multiple-choice questions with a single answer, which may not accurately gauge a model’s comprehension and can lead to superficial learning, i.e., the model will not look into other choices if the correct choice is straightforward. A more robust, detailed, and multi-scale dataset is necessary to effectively evaluate MLLMs under diverse conditions and scenarios. Current benchmarks mentioned above are evaluated with English context, while the rapid progression of Chinese MLLMs highlights the need for a Chinese multimodal benchmark with Chinese contents both in text and image and brings new challenges to the community.
In this paper, we introduce Multi, a novel benchmark named Multimodal Understanding Leaderboard with Text and Images, specifically designed to evaluate multimodal LLMs on cross-modal questions. Multi comprises 18,430 questions sourced from various educational and online materials, with most questions undergoing multiple rounds of human annotation for quality assurance. These questions cover a variety of scientific disciplines, including mathematics, physics, computer science, etc., and also pose significant challenges to intricate image reasoning. Multi serves as the first benchmark incorporating driving tests and administrative aptitude tests in China. The questions are crafted to test understanding and generation in various formats and complexity levels and are categorized into multiple-choice (with single or multiple correct answers), fill-in-the-blank, and open-ended questions.
To further challenge multimodal LLMs, we develop two subsets within Multi: Multi-Elite  consists of 500 carefully selected tough questions aiming to probe the limits of these models, and Multi-Extend  featuring 4,596 knowledge pieces tests the models’ capabilities of learning and knowledge transfer. These subsets offer deeper insights into the strengths and weaknesses of multimodal LLMs, fostering new research avenues. An example of Multi is shown in Figure 1, and more are presented in Appendix F.
We conduct comprehensive experiments on Multi using leading-edge multimodal and single-modality LLMs. Our findings reveal that multimodal LLMs still lag behind human performance in many aspects of Multi, highlighting challenges like cross-modal alignment, logical reasoning, mathematical computations, and image comprehension. Results show that the benchmark is challenging for current models, not to mention the Multi-Elite set where GPT-4V only gets a 14.0% score, and most of the other models get a score near random, indicating a large space for improvement.
In conclusion, We make the following contributions in this work:
  • We propose Multi, a substantial and challenging multimodal benchmark focusing on Chinese scientific questions, designed to evaluate multimodal LLMs.
  • We introduce Multi-Elite and Multi-Extend sets to test models’ bottleneck and in-context learning abilities, aiming for a more nuanced evaluation of multimodal LLMs.
  • We present detailed experiments with various state-of-the-art multimodal and single-modality LLMs on Multi, providing both qualitative and quantitative insights into their performance.
  • We make the Multi leaderboard, dataset, evaluation code, and the two subsets available to the research community, encouraging further participation and advancement in the field of multimodal LLMs.
@@ 2 Related Works
With advancements in aligning features across multiple modalities, like CLIP Radford et al. (2021) and ALBEF Li et al. (2021), recent studies have explored projecting vision features into the latent space of LLMs, aiming to enhance their capabilities of comprehending visual information. For example, BLIP-2 Li et al. (2023c) pioneers this approach by employing Q-Former to translate image features into text representations. Following this, LLaVA Liu et al. (2023b), MiniGPT-4 Zhu et al. (2023), and InstructBLIP Dai et al. (2023) have introduced visual instruction tuning to bolster the capability of MLLMs of following instructions. Our primary focus is on the proficiency of MLLMs in comprehending instructions in Chinese, which are divided into two main branches: open-source models, which typically build upon existing Chinese LLMs or are fine-tuned on Chinese instruction datasets, examples of which include Chinese-LLaVA LinkSoul-AI (2023), VisualGLM Du et al. (2022), VisCPM Hu et al. (2023), Qwen-VL Bai et al. (2023a), InternVL Zhang et al. (2023a), Yi-VL 01.ai (2023); and closed-source models, which are often highly powerful, multi-lingual systems such as GPT-4V(ision) OpenAI (2023b) and Gemini Team (2023). In this paper, we intend to evaluate these models across a range of scientific fields on the Multi benchmark, offering an extensive assessment and guidance for the onward trajectory of Chinese MLLMs.
@@
In assessing MLLMs, traditional methods primarily rely on established vision-language (VL) benchmark datasets. Renowned benchmarks such as VQA Goyal et al. (2017), OK-VQA Antol et al. (2015), GQA Hudson and Manning (2019), and MSCOCO Lin et al. (2014) are tailored to specific VL tasks like image captioning, open-domain visual question answering, and visual reasoning. While the evaluation based on standard benchmark datasets yields significant insights into MLLMs’ capabilities, these approaches may not entirely capture their comprehensive intelligence in real-world scenarios. Therefore, a diverse array of benchmarks has been developed to examine MLLMs on dealing with various tasks in real world. Benchmarks like LLaVA-Bench Liu et al. (2023b), MMBench Liu et al. (2023c), MM-VET Yu et al. (2023), TouchStone Bai et al. (2023b), MLLM-bench Ge et al. (2023), and SEED-BenchLi et al. (2023b, a), for instance, leverage GPT to evaluate the relevance and helpfulness of human-like long responses in the reality. POPE Li et al. (2023d) and HallusionBench Liu et al. (2023a) introduce various analytical criteria for the holistic evaluation of MLLMs’ hallucinations. Furthermore, M3Exam Zhang et al. (2023b), SciGraphQA Li and Tajbakhsh (2023), MathVista Lu et al. (2023), AGIEval Zhong et al. (2023), and MMMU Yue et al. (2023) consider MLLMs as experts to extend the evaluation scope by incorporating advanced perception and reasoning within domain-specific knowledge, for example, scientific questions and driving tests. The works most related to us are M3Exam, ScienceQA, SciEval Sun et al. (2023a) and C-Eval Huang et al. (2023). Our approach distinguishes itself by offering a broader spectrum of question types compared to the first two and supports a multimodal evaluation in contrast to the last two.
@@ 3 The Multi Benchmark
We propose Multi, a Multimodal Understanding Leaderboard with Text and Images, which can serve as a challenging and diverse benchmark for the MLLM community. The detailed statistics are provided in Appendix A.
@@ 3.1 Data Construction Process
Figure 3: The construction pipeline of Multi.
@@
We collect more than 2.7M raw data from the Internet, ranging from exams and quizzes from Chinese junior and senior schools and several society exams. We design an algorithm to pick out a proportion of the questions as the fundamental data of our benchmark. The selection is based on the questions’ text length, number of images, corresponding subjects, and knowledge pieces, to reach a higher diversity of questions and coverage of knowledge. The details are presented in Appendix D. We also collect questions from internal exams and practices of several top universities. After the selection, we obtain over 18K questions as the raw data.
@@
During the Data Annotation stage, an online platform facilitates annotators, primarily skilled undergraduates (involved in the work as authors), in tasks across format, content, label, and semantic levels. This includes converting content into markdown and , splitting sub-questions into individual ones, evaluating the difficulty and quality, and correcting errors for factual accuracy.
The Data Post-process stage employs strategies like formation, disambiguration, distillation, and transformation to enhance question difficulty and diversity, including modifying question formats and reducing assistance information.
Throughout these stages, we process 2.7 million questions in total and pick out 18,430, incorporating 23,320 scoring points, 7,658 images, and 4,595 knowledge pieces. Multi highlights a broad diversity in question types, including multiple-choice questions with both single and multiple answers, along with fill-in-the-blank and open-ended writing questions enriching the testing scenarios.  [^1] For the sake of simplifying writing, in the following paragraphs we may use abbreviations. We denote multiple-choice questions with a single answer as SA or Single Answer Choosing and those with multiple answers as MA or Multi Answer Choosing. We use FB for fill-in-the-blank questions and OP for open-ended writing questions. The stages during data processing and annotation significantly increase the diversity and difficulty of the dataset. For details of data processing and annotation, please refer to Appendix E.
@@ 3.2 The Multi-Elite Set
We select an additional set of 500 questions to create the advanced dataset. This set is comprised of objective questions, i.e. multiple-choice and fill-in-the-blank questions. The questions are averagely distributed in all of the subjects and education levels, evaluated as with high difficulty and quality by annotators, and with rich text and image content. The evaluation results presented in § 4 are also referred to in the selection, where the results of GPT-4VOpenAI (2023b) are given the most consideration.
@@ 3.3 The Multi-Extend Background Knowledge Dataset
External knowledge is crucial to provide critical information that assists in solving questions using the In-Context Learning (ICL) abilities. Some of the raw questions retrieved from the Internet have corresponding knowledge pieces attached. We also collect more knowledge pieces for uncovered questions with the assistance of LLMs and outer knowledge source (e.g. New Bing [^2] https://bing.com/new and Wikipedia [^3] https://wikipedia.org). We conduct annotations on these knowledge pieces to confirm the correctness of the content and present them in the Multi-Extend dataset. This dataset consists of about 4.6K knowledge pieces, designed to test the in-context learning abilities and knowledge transfer skills of models. This dataset provides comprehensive insights into the capabilities and limitations of multimodal LLMs, opening new pathways for research exploration.

### v3

@@ Abstract
The rapid development of multimodal large language models (MLLMs) raises the question of how they compare to human performance. While existing datasets often feature synthetic or overly simplistic tasks, some models have already surpassed human expert baselines. In this paper, we present Multi, a Chinese multimodal dataset derived from authentic examination questions. Comprising over 18,000 carefully selected and refined questions, Multi evaluates models using real-world examination standards, encompassing image-text comprehension, complex reasoning, and knowledge recall. Additionally, We also introduce Multi-Elite, a 500-question selected hard subset, and Multi-Extend with more than 4,500 external knowledge context pieces for testing in-context learning capabilities. Our evaluation highlights substantial room for MLLM advancement, with Qwen2-VL-72B achieving a 76.9% accuracy on Multi and 53.1% on Multi-Elite leading 25 evaluated models, compared to human expert baselines of 86.1% and 73.1%. Multi serves not only as a robust evaluation platform but also paves the way for the development of expert-level AI. Details and access are available at https://OpenDFM.github.io/MULTI-Benchmark.
@@
Zichen Zhu1and Yang Xu1and Lu Chen [^2] The corresponding authors are Lu Chen and Kai Yu.12and Jingkai Yang1and Yichuan Ma1 Yiming Sun1and Hailin Wen1and Jiaqi Liu1and Jinyu Cai1 Yingzi Ma1and Situo Zhang1and Zihan Zhao1and Liangtai Sun1and Kai Yu [^2] The corresponding authors are Lu Chen and Kai Yu.12 1 X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China 2 Suzhou Laboratory, Suzhou, China {JamesZhutheThird, xuyang0112, chenlusz, kai.yu}@sjtu.edu.cn
1 Introduction
The rapid development of large-scale language models (LLMs) Sun et al. (2023); Chen et al. (2022); qwe (2024); Cai et al. (2024); OpenAI (2022, 2023a); Team (2023); Touvron et al. (2023) has brought about remarkable progress in natural language processing (NLP) and related fields. However, human communication and understanding extend far beyond just language, encompassing various other forms of information such as images, mathematical formulas, graphs, tables, and diagrams. These visual modalities play an essential role in fields like science and engineering, where they are crucial for conveying complex ideas. Consequently, there has been increasing interest in developing Multimodal Large Language Models (MLLMs) Du et al. (2022); Hu et al. (2023); LinkSoul-AI (2023); Bai et al. (2023a); 01.ai (2023); Chen et al. (2024); Zhe et al. (2024); Yao et al. (2024); Wang et al. (2024); OpenAI (2023b); Team (2023); OpenAI (2024a, b); Anthropic , that can process and generate across different modalities, including visual ones, and perform tasks that require cross-modal reasoning. All these rapid advancements raise the question of how a model compares to a real student.
Figure 1: Multi is a large-scale Chinese multimodal benchmark across multiple domains and aspects.
Evaluating MLLMs presents unique challenges, as current benchmarks Lu et al. (2022); Li et al. (2024a) are either too narrow in scope, focusing on natural scene images, or too simplistic to effectively assess the capabilities of modern models. Many existing scientific benchmarks Sun et al. (2024); Huang et al. (2024) rely on multiple-choice questions with a single correct answer, which may not adequately test a model’s comprehension or reasoning skills. As a result, models can sometimes arrive at superficial solutions without fully understanding the problem. A more robust, detailed, and multi-scale dataset is necessary to evaluate MLLMs in a more comprehensive manner under diverse real-world conditions. Additionally, while most existing benchmarks are in English, the rapid progress of Chinese multimodal models has highlighted the need for a Chinese multimodal benchmark that includes both Chinese text and visual content, offering unique challenges to the field.
In this paper, we introduce Multi, a new multimodal benchmark for evaluating LLMs on cross-modal understanding tasks. Our primary goal was to evaluate MLLMs across a broad range of tasks that a typical Chinese student would encounter throughout their academic progression. The Multi benchmark, which stands for Multimodal Understanding Leaderboard with Text and Images, consists of 18,430 carefully curated questions derived from real-world educational and online sources. These questions span multiple scientific domains, including mathematics, physics, and computer science, and feature a variety of question formats, such as multiple-choice, fill-in-the-blank, and open-ended questions. Importantly, Multi also includes driving and administrative aptitude tests, which are significant in the Chinese context and offer further challenges in cross-modal reasoning and generation.
To push the boundaries of MLLM evaluation, we also introduce two specialized subsets within Multi: Multi-Elite, which includes 500 tough questions designed to test the limits of current MLLMs, and Multi-Extend, a set of 4,596 external knowledge pieces that assess models’ abilities in in-context learning and knowledge transfer. These subsets provide deeper insights into the strengths and weaknesses of MLLMs, driving further research and model improvement. Sample questions from Multi are presented in Figure 2, and additional details are available in E and F.
@@ 1 Introduction
We conduct extensive experiments on Multi, including both multimodal and single-modality LLMs, including 14 open-sourced and 11 close-source models. Our results show that while multimodal LLMs have made significant progress, they still lag behind human expert performance on many aspects of the Multi benchmark. Key challenges remain in cross-modal alignment, logical reasoning, mathematical computation, and image comprehension. In particular, the Multi-Elite subset proves especially challenging, with the best-performing model, Qwen2-VL-72B, achieving only a 50.0% accuracy, and many models performing near random levels. This highlights a substantial space for improvement in current MLLMs.
In summary, the contributions of this paper are as follows:
  • We introduce Multi, a comprehensive and challenging multimodal benchmark designed to evaluate Chinese scientific knowledge tasks for MLLMs.
  • We present the Multi-Elite and Multi-Extend subsets to evaluate models’ reasoning capabilities and in-context learning abilities, providing a more nuanced assessment of MLLMs.
  • We conduct detailed experiments with 25 state-of-the-art multimodal and single-modality models, offering both qualitative and quantitative insights into their performance on Multi. We also conducted a bunch of ablation studies to draw further conclusions.
  • We conduct human evaluations on 1k questions and give error analysis on 1.3k responses from models.
@@ 2 Related Works
With advancements in aligning features across multiple modalities, like CLIP Radford et al. (2021) and ALBEF Li et al. (2021), recent studies have explored projecting vision features into the latent space of LLMs, aiming to enhance their capabilities of comprehending visual information. For example, BLIP-2 Li et al. (2023a) pioneers this approach by employing Q-Former to translate image features into text representations. Following this, LLaVA Liu et al. (2024), MiniGPT-4 Zhu et al. (2024), and InstructBLIP Dai et al. (2024) have introduced visual instruction tuning to bolster the capability of MLLMs of following instructions. Our primary focus is on the proficiency of MLLMs in comprehending instructions in Chinese, which are divided into two main branches: open-source models, which typically build upon existing Chinese LLMs or are fine-tuned on Chinese instruction datasets, examples of which include Chinese-LLaVA LinkSoul-AI (2023), VisualGLM Du et al. (2022), VisCPM Hu et al. (2023), Qwen-VL Bai et al. (2023a), InternVL Zhang et al. (2023), Yi-VL 01.ai (2023); and closed-source models, which are often highly powerful, multi-lingual systems such as GPT-4V(ision) OpenAI (2023b) and Gemini Team (2023). In this paper, we intend to evaluate these models across a range of scientific fields on the Multi benchmark, offering an extensive assessment and guidance for the onward trajectory of Chinese MLLMs.
@@
In assessing MLLMs, traditional methods primarily rely on established vision-language (VL) benchmark datasets. Renowned benchmarks such as VQA Goyal et al. (2017), OK-VQA Antol et al. (2015), GQA Hudson and Manning (2019), and MSCOCO Lin et al. (2014) are tailored to specific VL tasks like image captioning, open-domain visual question answering, and visual reasoning. While the evaluation based on standard benchmark datasets yields significant insights into MLLMs’ capabilities, these approaches may not entirely capture their comprehensive intelligence in real-world scenarios. Therefore, a diverse array of benchmarks has been developed to examine MLLMs in dealing with various tasks in the real world. Benchmarks like LLaVA-Bench Liu et al. (2024), MMBench Liu et al. (2025), MM-VET Yu et al. (2024), TouchStone Bai et al. (2023b), MLLM-bench Ge et al. (2023), and SEED-BenchLi et al. (2024a, b), for instance, leverage GPT to evaluate the relevance and helpfulness of human-like long responses in the reality. POPE Li et al. (2023b) and HallusionBench Guan et al. (2024) introduce various analytical criteria for the holistic evaluation of MLLMs’ hallucinations. Furthermore, M3Exam Zhang et al. (2024b), SciGraphQA Li and Tajbakhsh (2023), MathVista Lu et al. (2024), AGIEval Zhong et al. (2024), and MMMU Yue et al. (2024) consider MLLMs as experts to extend the evaluation scope by incorporating advanced perception and reasoning within domain-specific knowledge, for example, scientific questions and driving tests. The works most related to us are M3Exam, ScienceQA, SciEval Sun et al. (2024) and C-Eval Huang et al. (2024). Our approach distinguishes itself by offering a broader spectrum of question types compared to the first two and supports a multimodal evaluation in contrast to the last two.
@@ 3 The Multi Benchmark
We propose Multi, a Multimodal Understanding Leaderboard with Text and Images, which can serve as a challenging and diverse benchmark for the MLLM community. In § 3.1, we provide an overall introduction to the data construction process, with more details provided in E. The detailed statistics are provided in § 3.2. In § 3.3 and § 3.4 we introduce Multi-Elite and Multi-Extend. We also give a comparison between Multi and other knowledge-based benchmarks in § 3.5
@@ 3.1 Data Construction Process
Figure 3: The construction pipeline of Multi
@@
We collect more than 2.7M raw data from the Internet, ranging from exams and quizzes from Chinese junior and senior schools and several society exams. We design an algorithm to pick out a proportion of the questions as the fundamental data of our benchmark. The selection is based on the questions’ text length, number of images, corresponding subjects, and knowledge pieces, to reach a higher diversity of questions and coverage of knowledge. The details are presented in D. We also collect questions from internal exams and practices of several top universities. After the selection, we obtain over 18K questions as the raw data.
@@
During the Data Annotation stage, an online platform facilitates annotators, primarily skilled undergraduates (involved in the work as authors), in tasks across the format, content, label, and semantic levels. This includes converting content into markdown and , splitting sub-questions into individual ones, evaluating the difficulty and quality, and correcting errors for factual accuracy.
The Data Post-process stage employs strategies like formation, disambiguation, distillation, and transformation to enhance question difficulty and diversity, including modifying question formats and reducing assistance information.
Throughout these stages, we process 2.7 million questions in total and pick out 18,430, incorporating 23,320 scoring points, 7,658 images, and 4,595 knowledge pieces. Multi highlights a broad diversity in question types, including multiple-choice questions with both single and multiple answers, along with fill-in-the-blank and open-ended writing questions enriching the testing scenarios.  [^1] For the sake of simplifying writing, in the following paragraphs we may use abbreviations. We denote multiple-choice questions with a single answer as SA or Single Answer Choosing and those with multiple answers as MA or Multi Answer Choosing. We use FB for fill-in-the-blank questions and OP for open-ended writing questions. The stages during data processing and annotation significantly increase the diversity and difficulty of the dataset. For details of data processing and annotation, please refer to E.
3.2 Statistics
We provide detailed statistics in Table 1. We also provide the composition of education level and subjects as shown in Figure 4.
Statistics Number Points
Total Problems 17250 -
Total Questions 18429 -
Total Points 22926 -
Total Images 7734 -
Total Knowledge 4595 -
Multiple Choices 16095(87.33%) 19906(86.83%)
- Single Answer 13949(75.69%) 13952(60.86%)
- Multiple Answers 2143(11.63%) 5954(25.98%)
Fill in the Blank 1439(7.81%) 2223(9.69%)
Open Ended Writing 797(4.33%) 797(3.48%)
Others 98(0.53%) -
Question with Images 7521(40.81%) 8836(38.54%)
- Single Image 7279(39.50%) 8563(37.35%)
- Multiple Images 242(1.21%) 273(1.19%)
- Image in Choice 1179(6.40%) 1181(5.06%)
Question with Explanations 10565(57.34%) 13186(57.52%)
Question with Knowledge 9048(49.10%) 12919(55.35%)
Table 1: The statistic overview of Multi. Since Multi contains questions worth multiple points, we provide distribution statistics of both measurements. The statistics are based on v1.3.0-20241203.
Figure 4: The question distribution by education level and subjects
Type #choices #A #B #C #D #E,F,G...
SA 2 1813 1370 - - -
3 273 288 265 - -
4 2192 2636 2702 2379 -
5 3 6 4 4 0
[…]
Total 2-13 5747 5871 4490 3694 92
Table 2: The answer distribution for multiple-choice questions
In addition to multiple-choice questions, our benchmark also includes a substantial number of fill-in-the-blank and open-ended questions, creating a diverse and comprehensive range of testing scenarios. Moreover, we have incorporated unique open-response questions that require creative answers, such as drawings. It is important to note that these open-response questions are not included in our formal evaluation and scoring procedures, for they are primarily proposed for qualitative research and development in the field of MLLMs. Our benchmark is carefully designed to thoroughly assess and enhance the ability of MLLMs to process and respond to various question types, resembling real-world scenarios.
@@ 3.3 The Multi-Elite Set
We have selected 500 objective questions to create a hard subset. These multiple-choice and fill-in-the-blank questions are more evenly distributed across various educational levels and subjects and have been evaluated by annotators as high in difficulty and quality, featuring detailed content and complex reasoning requirements. In the selection process, the performance of several early-released models during earlier evaluation is also taken into consideration. We especially focused on poorly answered questions by GPT-4VOpenAI (2023b), the state-of-the-art model at the time. As a result, these models may get lower scores compared to the random baseline in Table 7. Since a complete run on Multi could be costly for some models, we encourage researchers to run their experiments on these 500 high-quality questions.
@@ 3.4 The Multi-Extend Background Knowledge Dataset
External knowledge is crucial to provide critical information that assists in solving questions using the In-Context Learning (ICL) abilities. Some of the raw questions retrieved from the Internet have corresponding knowledge pieces attached. We also collect more knowledge pieces for uncovered questions with the assistance of LLMs and outer knowledge sources (e.g. New Bing [^2] https://bing.com/new and Wikipedia [^3] https://wikipedia.org). We conduct annotations on these knowledge pieces to confirm the correctness of the content and present them in the Multi-Extend dataset. This dataset consists of about 4.6K knowledge pieces, designed to test the in-context learning abilities and knowledge transfer skills of models. This dataset provides comprehensive insights into the capabilities and limitations of multimodal LLMs, opening new pathways for research exploration.